how many molecules of carbon dioxide are produced when 67.9g of dicarbon hexahydride are reacting?


how many liters of oxygen are needed to react when 8.23 x 10^26 molecules of water are produced at STP? 

 Please show all work for both questions


1
Expert's answer
2021-02-28T06:18:34-0500

2C2H6 + 7O2 = 4CO2 + 6H2O

n(C2H6) = 67.9g/30g/mol = 2.26 mol

n(CO2) = 2*2.26 = 4.52 mol

N(CO2) = 4.52*6.02*10^23 = 27.2*10^23 molecules


2H2 + O2 = 2H2O

n(H2O) = 8.23*10^26/6.02*10^23 = 1.37*10^3 moles

n(O2) = 1.37*10^3/2 = 6.85*10^2 moles

V(O2) = 6.85*10^2mol*22.4L/mol = 15344 L